OpenStudy (anonymous):

For the simple harmonic motion equation 5sin((π/4)t), what is the period?

OpenStudy (anonymous):

period is \[\frac{2\pi}{b}\] an in this case \[b=\frac{\pi}{4}\]

OpenStudy (anonymous):

get \[2\pi\times \frac{4}{\pi}=8\]